# Continuous optimization by quantum adaptive distribution search

### v1

@@ IV.2 Performance estimation by classical algorithms
In the next experiment, we examine the performance of our method in higher dimensional settings. Since the size of the state vector in quantum simulators increases exponentially as the dimension increases, conducting experiments on a quantum simulator becomes difficult in more than 4 dimensions when $\tau=8$. To measure performance in higher dimensions, we replace the quantum sampling in QuADS and GAS with classical sampling and estimate the number of oracle calls required for the entire optimization in the case of quantum sampling from the optimization process of the corresponding classical algorithm.
@@
where $p$ represents the probability of obtaining a correct sample in the classical sampling. When we set the number of rotations $r$ in amplitude amplification to $N_{\mathrm{opt}}(p)$, and for $\theta$ satisfying $\sin^{2}(\theta)=p$, we find that $\sin^{2}((2N_{\mathrm{opt}}(p)+1)\theta)=1$. This result demonstrates that $N_{\mathrm{opt}}(p)$ indeed represents the optimal number of rotations. Thus, $N_{\mathrm{opt}}(p)$ serves as a lower bound of the expected oracle call count within each amplitude amplification step of GAS or QuADS. We estimate $p$ for each step by classical sampling; if obtaining $M_{i}$ samples under the threshold requires $K_{i}$ function evaluations at $i$th iteration of the corresponding classical algorithm, we estimate probability $p_{i}$ as $\tilde{p_{i}}=M_{i}/K_{i}$. We then estimate a lower bound of $o_{\rm local}$ ($o_{\rm global}$) by an average of $\sum_{i}N_{\mathrm{opt}}(\tilde{p_{i}})$ using the optimization trials that achieved local (global) convergence. Then, a lower bound to $o_{\rm total}$, which we denote as $\tilde{o}_{\rm lower}$, is estimated through Eq. (3) using the corresponding estimated lower bounds for $o_{\rm local}$ and $o_{\rm global}$. Finally, $o_{\rm total}$ can be inferred from $\tilde{o}_{\rm lower}$, by multiplying a certain constant coefficient. This approach is valid because both $o_{\rm total}$ and $\tilde{o}_{\rm lower}$ scale in a similar manner with respect to $D$. We conducted 100 individual optimizations to estimate $\tilde{o}_{\rm lower}$ for each of the rastrigin, schwefel, styblinski tang, ackley, alpine01, and griewank functions up to feasible dimensions in classical algorithms. We choose the functions because QuADS surpass GAS and CMA-ES greatly for the former two, while for the latter functions, differences between CMA-ES and QuADS are small (see Fig. 3).
@@
First, we checked the consistency of using $\tilde{o}_{\rm lower}$ to estimate $o_{\rm total}$. In Fig. 5, we plot $\tilde{o}_{\rm lower}$ against $o_{\rm total}$ of corresponding functions with the dimensions which are both feasible in the quantum and classical algorithm. It shows $o_{\rm total}$ is correctly lower bounded by $\tilde{o}_{\rm lower}$. Furthermore, it can be experimentally observed that $o_{\rm total}$ can be closely approximated by $o_{\rm total}=2\tilde{o}_{\rm lower}$. Thus, we use $\tilde{o}_{\rm total}:=2\tilde{o}_{\rm lower}$ as an estimator for $o_{\rm total}$ of GAS and QuADS in high dimensional setting.
@@
Figure 5: The consistency verification between classical estimation and quantum simulation regarding the expected number of oracle calls across the functions in Fig. 6. The dotted line represents $o_{\rm total}=\tilde{o}_{\rm lower}$. The classical estimation results ($\tilde{o}_{\rm lower}$) serve as a lower bound to the quantum simulation results ($o_{\rm total}$), which is consistent with theoretical predictions. The results of the quantum simulations are approximately twice that of the classical estimations, as indicated by the solid line $o_{\rm total}=2\tilde{o}_{\rm lower}$ for both QuADS and GAS.
@@ V Conclusion
This paper proposes a method for quantum continuous optimization in a quantum computer. In our method, we enhance GAS by integrating the efficient search capabilities of CMA-ES. Our method can efficiently utilize function structure in the search process through the updating rule of CMA-ES. Our numerical experiment shows our method surpasses both CMA-ES and GAS in terms of expected oracle call count for finding global optimum. QuADS has a smaller increase in oracle call count with respect to dimension, and this observation suggests our method’s significant potential in addressing high-dimensional problems commonly encountered in practical applications.
We list possible future studies in the following. First, although we neglected the cost for the initial state preparation, investigating the overall performance of QuADS including this cost is needed. However, we expect that the state preparation cost can be negligible for complex objective functions $f(x)$. Second, we can extend our approach to discrete variable cases. Utilizing distribution in discrete optimization is generally referred to as “estimation of distribution algorithm” [23]. Third, for the common challenge of continuous optimization on a quantum computer, we need to find practical oracles that are efficient in real time. The conditions for such an oracle would be multimodal functions which are difficult to optimize on a classical computer. Also, oracles that are more efficient when computed on a quantum computer, like the energy of materials, are attractive candidates because we can naturally incorporate our method.

### v2

@@ IV.2 Performance estimation by classical algorithms
Next, we examine the performance of our method in higher dimensional settings. As the dimensionality increases, the number of local optima grows exponentially, significantly complicating the optimization problem. Consequently, this leads to a reduction in $p$, enhancing the effectiveness of the quadratic acceleration inherent in quantum search. Therefore, it is anticipated that QuADS becomes increasingly advantageous under these conditions. Since the size of the state vector in quantum simulators increases exponentially as the dimension increases, conducting simulations on a quantum simulator becomes difficult in more than 4 dimensions when $\tau=8$. To measure performance in higher dimensions, we replace the quantum sampling in QuADS and GAS with classical sampling and estimate the number of oracle calls required for the entire optimization in the case of quantum sampling from the optimization process of the corresponding classical algorithm.
@@
where $p$ represents the probability of obtaining a correct sample in the classical sampling. When we set the number of rotations $r$ in amplitude amplification to $N_{\mathrm{opt}}(p)$, and for $\theta$ satisfying $\sin^{2}(\theta)=p$, we find that $\sin^{2}((2N_{\mathrm{opt}}(p)+1)\theta)=1$. This result demonstrates that $N_{\mathrm{opt}}(p)$ indeed represents the optimal number of rotations. Thus, $N_{\mathrm{opt}}(p)$ serves as a lower bound of the expected oracle call count within each amplitude amplification step of GAS or QuADS. We estimate $p$ for each step by classical sampling; if obtaining $M_{i}$ samples under the threshold requires $K_{i}$ function evaluations at $i$th iteration of the corresponding classical algorithm, we estimate probability $p_{i}$ as $\tilde{p_{i}}=M_{i}/K_{i}$. We then estimate a lower bound of $o_{\rm local}$ ($o_{\rm global}$) by an average of $\sum_{i}N_{\mathrm{opt}}(\tilde{p_{i}})$ using the optimization trials that achieved local (global) convergence. Then, a lower bound to $o_{\rm total}$, which we denote as $\tilde{o}_{\rm lower}$, is estimated through Eq. (2) using the corresponding estimated lower bounds for $o_{\rm local}$ and $o_{\rm global}$. Finally, $o_{\rm total}$ can be inferred from $\tilde{o}_{\rm lower}$, by multiplying a certain constant coefficient. This approach is valid because both $o_{\rm total}$ and $\tilde{o}_{\rm lower}$ scale in a similar manner with respect to $D$. We conducted 100 individual optimizations to estimate $\tilde{o}_{\rm lower}$ for each of the rastrigin, schwefel, styblinski tang, ackley, alpine01, and griewank functions up to feasible dimensions in classical algorithms. We choose the functions because QuADS surpasses GAS and CMA-ES greatly for the former two, while for the latter functions, differences between CMA-ES and QuADS are small (see Fig. 3).
@@
First, we checked the consistency of using $\tilde{o}_{\rm lower}$ to estimate $o_{\rm total}$. Figure 5 shows the relationship between $\tilde{o}_{\rm{lower}}$ and $o_{\rm{total}}$. Each data point in the plot corresponds to a specific function and dimension for which both $\tilde{o}_{\rm{lower}}$ and $o_{\rm{total}}$ can be computed. It shows $o_{\rm total}$ is correctly lower bounded by $\tilde{o}_{\rm lower}$. Furthermore, it can be observed from simulation that $o_{\rm total}$ can be closely approximated by $o_{\rm total}=2.3\tilde{o}_{\rm lower}$. Thus, we use $\tilde{o}_{\rm total}:=2.3\tilde{o}_{\rm lower}$ as an estimator for $o_{\rm total}$ of GAS and QuADS in high dimensional setting. More detail analysis about consistency of this approximation in high dimension is shown in Appendix D.
@@
Figure 5: The consistency verification between classical estimation and quantum simulation regarding the expected number of oracle calls across the functions in Fig. 6. The dotted line represents $o_{\rm total}=\tilde{o}_{\rm lower}$. The classical estimation results ($\tilde{o}_{\rm lower}$) serve as a lower bound to the quantum simulation results ($o_{\rm total}$), which is consistent with theoretical predictions. The results of the quantum simulations are approximately twice that of the classical estimations, as indicated by the solid line $o_{\rm total}=2.3\tilde{o}_{\rm lower}$ for both QuADS and GAS.
@@ V Conclusion
This paper proposes a method for quantum continuous optimization in a quantum computer. In our method, we enhance GAS by integrating the efficient search capabilities of CMA-ES. Our method can efficiently utilize function structure in the search process through the updating rule of CMA-ES. Our numerical simulation shows our method surpasses both CMA-ES and GAS in terms of expected oracle call count for finding global optimum. QuADS has a smaller increase in oracle call count with respect to dimension, and this observation suggests our method’s significant potential in addressing high-dimensional problems commonly encountered in practical applications.
We list possible future studies in the following. First, although we neglected the cost for the initial state preparation, investigating the overall performance of QuADS including this cost is needed. However, we expect that the state preparation cost can be negligible for complex objective functions $f(x)$. Second, we can extend our approach to discrete variable cases. Utilizing distribution in discrete optimization is generally referred to as “estimation of distribution algorithm” [26]. Third, for the common challenge of continuous optimization on a quantum computer, we need to find practical oracles that are efficient in real time. The conditions for such an oracle would be multimodal functions which are difficult to optimize on a classical computer. Also, oracles that are more efficient when computed on a quantum computer, like the energy of materials, are attractive candidates.
@@ VI Acknowledgement
Appendix A Hyperparameters for CMA-ES Update
Variable
$w_{i}$ $\frac{\log(K+1/2)-\log(i)}{\sum_{i=1}^{K}\left\|\log(K+1/2)-\log(i)\right\|}$
$\mu_{\text{eff}}$ $\frac{1}{\sum_{i=1}^{K}w_{i}^{2}}$
$c_{1}$ $\frac{2}{(D+1.3)^{2}+\mu_{\text{eff}}}$
$c_{\sigma}$ $\frac{\mu_{\text{eff}}+2}{D+\mu_{\text{eff}}+5}$
$c_{\mu}$ $\min\left(1-c_{1},2\left(\mu_{\text{eff}}-2+\frac{1}{\mu_{\text{eff}}}\right)\right)$
$d_{\sigma}$ $\min\left(1-c_{1},\frac{2(\mu_{\text{eff}}-2+\frac{1}{\mu_{\text{eff}}})}{(D+2)^{2}+\mu_{\text{eff}}}\right)$
Table 2: Hyperparameters setting for CMA-ES update
In this section, we describe hyperparameters used in Section IV. For CMA-ES update (Algorithm 4) in CMA-ES and QuADS, we employ the default hyperparameters used in [14]. The hyperparameters for CMA-Update are enumerated in Table 2.
Appendix B Particle swarm optimization and basin hopping optimizer
PSO [23, 24] is a swarm-based optimizer that iteratively improves candidate solutions through communication among the swarm. PSO simulates the social behavior of birds within a flock. The position of each bird in the search space represents a potential solution, and its flight is guided by its own experience as well as the experience of other birds. The update equations for the velocity and position of each particle can be written as:
$v_{i}^{(t+1)}$ $=wv_{i}^{(t)}+c_{1}r_{1}(p_{i}-x_{i}^{(t)})+c_{2}r_{2}(p_{g}-x_{i}^{(t)})$
$x_{i}^{(t+1)}$ $=x_{i}^{(t)}+v_{i}^{(t+1)}$
where $v_{i}^{(t)}$ is the velocity of particle $i$ at iteration $t$, $x_{i}^{(t)}$ is the current position of particle $i$, $p_{i}$ is the best known position of particle $i$, $p_{g}$ is the global best known position among all particles, $w$ is the inertia weight, $c_{1}$ and $c_{2}$ are the personal and global coefficients, respectively, and $r_{1}$ and $r_{2}$ are random numbers between 0 and 1. For the implementation of PSO, the PySwarm library [27] was utilized, with the hyperparameters set as $w=0.9$, $c_{1}=0.5$, and $c_{2}=0.3$.
Basin-hopping is a global optimization algorithm that enhances a local optimization method through the integration of stochastic perturbations to the coordinates, enabling the algorithm to escape local minima. In our implementation, the basinhopping module from SciPy optimization library[28] was employed. Due to the default configuration does not handle bounds on the optimization function, we modified the random perturbation process to ensure it does not exceed the defined bounds. If a perturbation results in exceeding the bounds, the solution is adjusted back to the nearest boundary. As a subroutine of the basinhopping algorithm, the L-BFGS-B method [29] was adopted.
Appendix C Test functions
(a) rastrigin function.
(b) ackley function.
(c) alpine01 function.
(d) alpine02 function.
(e) deflected corrugated spring function.
(f) griewank function.
(g) schwefel function.
(h) styblinski function.
(i) wavy function.
Figure 7: The landscapes of two dimensional test functions in Table 1.
Figure 8: The landscape of one dimensional griewank function. The red dots indicate the minimum point of the griewank function.
The plots of test functions in Table 1 are shown in Figure 7 for the two-dimensional case. Although the griewank function looks similar to a quadratic function in Figure 7, it actually exhibits significant fluctuations. A more detailed view of griewank function at a smaller scale is provided in Figure 8.
Appendix D Consistency of approximation of the number of oracle calls with classical simulations
In Section IV.2, we estimate oracle counts of quantum computation in high dimension from the number of function calls obtained from classical computations using the equation $o_{\rm total}=2.3\tilde{o}_{\rm lower}$. Does the relation hold true for higher dimensions as well? This section aims to demonstrate the validity of this transformation in higher dimension. Specifically, we investigated the ratio of the expected number of oracle calls required for amplitude amplification to $N_{opt}(p)$ when $\lambda=5/4$ for each $p$.
First, let’s derive the expression for the expected number of oracle calls in amplitude amplification when $p$ is fixed. The maximum rotation count $n_{k}$ in the $k$-th trial is given by $n_{k}:=\left\lfloor\lambda^{k-1}\right\rfloor$ and the acceptance probability $a_{r}$ after performing $r$ amplitude amplifications is
$a_{r}:=\sin^{2}\left(\left(2r+1\right)\arcsin{\sqrt{p}}\right).$ ) .
In the $k$-th trial, we uniformly select $r$ from integers $0,\ldots,\left\lfloor\lambda^{k-1}\right\rfloor$ and perform $r$ amplitude amplifications. Thus, the probability of rejection $b_{k}$ in the $k$-th trial is expressed as
$b_{k}:=1-\frac{1}{n_{k}+1}\sum_{r=0}^{n_{k}}a_{r}.$
Therefore, the expected total number of oracle calls $S$ required until acceptance is computed as follows:
$S(p)$ $=\lim_{K\to\infty}\sum_{k=1}^{K}\sum_{r=0}^{n_{k}}\frac{1}{n_{k}+1}r\left(\prod_{j=1}^{k-1}b_{j}\right)a_{r}$
$=\lim_{K\to\infty}\sum_{k=1}^{K}\frac{1}{n_{k}+1}\left(\prod_{j=1}^{k-1}b_{j}\right)\left(\sum_{r=0}^{n_{k}}ra_{r}\right).$
We compute this sum of infinite series by finite approximation with a sufficiently large $K$. Figure 9 shows the ratio $S(p)/N_{opt}(p)$ for each $p$. This ratio converges around $2.3$ when $p$ is small, which is consistent with Fig. 5.
Figure 9: $S(p)/N_{opt}(p)$ for each $p$.